Given demand function: Q = 20 + m/20p
m=480 p=10, calculate the demand for the product given p=10
B) calculate demand for product when o falls to 4
C) calculate substitution effect when p falls to 4
D) calculate the income effect where price is held at 4
1
Expert's answer
2018-04-16T08:56:13-0400
Q = 20 + m/20p, m = 480, p = 10.
A) The demand for the product given p = 10 is: Q1 = 20 + 480/(20*10) = 22.4 units.
B) The demand for product when p falls to 4 is: Q2 = 20 + 480/(20*4) = 26 units.
C) Total effect = 26 - 22.4 = 3.6,
change in m = Q1(P2 - P1) = 22.4*(4 - 10) = -134.4,
m' = 480 - 134.4 = 345.6,
Q3 = 20 + 345.6/(20*4) = 24.32,
The substitution effect when p falls to 4 is: Q3 - Q1 = 24.32 - 22.4 = 0.92 units,
D) the income effect where price is held at 4 is: Q2 - Q3 = 26 - 24.32 = 1.68 units.